OpenStudy (anonymous):

The number of cubic feet in the volume of a cube is the same as the number of square feet in it's surface area. The length of the edge of the cube in feet is.....

OpenStudy (anonymous):

V=s^3 SA=6s^2 s^3=6s^2

OpenStudy (anonymous):

dividing each side by s^2 gives us s=6